OpenStudy (anonymous):

what percent of 74 is 481 ? if necessary, round your answer to the nearest tenth of a percent.

OpenStudy (anonymous):

= 100%(481/74) = 100%([481/37]/[74/37]) = 100%(13/2) = 1,300/2% or 650%

OpenStudy (anonymous):

Can i have medal

OpenStudy (anonymous):

@Ineedhelpthankyou you copy pig

OpenStudy (anonymous):

YAY

OpenStudy (anonymous):

you copied from yahoo answers

OpenStudy (anonymous):

here is proof http://answers.yahoo.com/question/index?qid=20091003192431AAUAcGE

OpenStudy (anonymous):

answer it by your own

OpenStudy (goldphenoix):

Even though he copied it in answer.yahoo, he "showed" the answer with explaniation. Now that is morally wrong in my opinion, but it's still the proper way of explaining.

OpenStudy (anonymous):

Ha thx gold

OpenStudy (anonymous):

shuh!!

OpenStudy (anonymous):

ya i did copy but i got the answer

OpenStudy (anonymous):

If you ever need help im me black beauty